# CCaaLF: Concurrency Control as a Learnable Function

### v3

@@ 9             $S\leftarrow$ all operations from conflict transactions
14                        bar $\leftarrow$ $w_{T^{\prime}.type}$ $S\leftarrow S\cup\{\text{1-st}\text{ to }\text{bar-th}\text{ operations in }T^{\prime}\}$
15
16            else if  EarlyValidation$(T)$ = Abort  then
@@ 23      $result\leftarrow$ SafeExecute($T,op$)
26             if  EarlyValidation$(T)$ = Abort  then
@@ 4.5.1. Execution
CCaaLF does not guarantee all conflicts are detected and resolved during execution, similar to OCC. A transaction $T$ has the following fields:
@@
For each operation $op$, we maintain the following information:
@@
  • $op.key$: the primary key of the data object accessed by $op$.
@@
  • $op.priority$: the priority of $op$ when resolving its conflicts with others operations.
@@
Algorithm 1 describes how a transaction operation is executed with learned actions. It performs early validation to proactively abort transactions that will fail later (line  1). The algorithm also enforces early validation before exposing a write (line 1), where the failure of a transaction $T$ could lead to the cascading abort of all transactions depending on it. By performing early validation (Algorithm 2), it reduces the impact of cascading aborts. The algorithm does not perform early validation for “detect all” because doing so incurs additional overhead and can restrict the search space. For example, performing early validation in “detect all” would prevent CCaaLF from replicating the behavior of 2PL, which operates without the overhead of early validation checks. Furthermore, our algorithm distinguishes between stored procedure mode and interactive transaction mode. In the former, all the operations of a transaction are known prior to transaction execution. In the latter, they are not known before transaction execution.
@@
For interactive transactions, the algorithm disallows certain actions that could compromise correctness due to the transaction’s logic. Such actions include dirty reads and partial retry, because the behavior of the conflicting transaction is unknown, and because the user has already seen the outputs of some operations. Before executing $op$, the algorithm calls $GetCC(T,op)$ to get actions $\mathbf{{a}}$ (line 1). Next, it executes conflict detection and stores the conflicts in $S$. This step is skipped if ${\mathbf{{a}}}_{d}$ is no detection, and the algorithm enforces $op$ to read clean data (lines 11). If ${\mathbf{{a}}}_{d}$ is detect critical, it performs early validation to check if previously read data have changed (line 1) and aborts $T$ if the validation fails. The conflicts with lower priorities than that of $op$ are removed from $S$. The remaining are resolved with the given timeout ${\mathbf{{a}}}_{t}$ (lines 11). If not all the conflicts in $S$ are resolved within ${\mathbf{{a}}}_{t}$, $T$ is aborted (line 1). Otherwise, $T$ proceeds to read the latest committed data (line 1), or to write to its write set $T.ws$ (line 1).
@@
For stored procedure transactions, the algorithm enables more flexibility by allowing dirty reads. It uses $w_{1},\dots,w_{n}$ for detecting critical conflicts before dirty reads, and $expose$ for deciding whether to make uncommitted write visible to other transactions. Each data object has a version chain containing dirty data. The execution when ${\mathbf{{a}}}_{d}$ is no detection or detect all is the same as that of interactive transactions. When ${\mathbf{{a}}}_{d}$ is detect critical, it detects critical conflicts between $T$ and its dependent transactions in $T.dep$ (lines  11). In particular, for every $T^{\prime}\in T.dep$, it adds the first $w_{T^{\prime}.type}$ operations of $T^{\prime}$ to $S$. Once all the conflicts in $S$ are resolved, it lets $op$ read the data at the tail of the corresponding version chain, and updates $T.dep$ to track the read-write dependency (lines 11). After executing $op$ (line 1), if $expose=1$, it performs early validation and aborts $T$ if the validation fails. If the validation succeeds, it exposes the uncommitted data in $T.ws$ by appending them to the corresponding version chains.
1 Function EarlyValidation($T$):
@@ 9 Function Commit($T$):
18             if $op.tx\neq op^{\prime}.tx$ then
19                   return Abort
20
@@
23             Append $op$ to tuple($op.key$)’s version chain as clean data
@@ 4.5.2. Validation and commit
Algorithm 2 shows how the algorithm performs an OCC-like four-step validation at commit time. First, it waits for all dependent transactions in $T.dep$ to complete, and aborts if any dependent transaction aborts (lines 22). Second, it locks the private write set to avoid concurrent read of the data in the set (line 2). Third, it verifies that data in the read set are the same as the latest committed versions (lines 22). It aborts if the verification fails. Fourth, it removes dirty data from $T$ on the corresponding version chains and unlocks the write set (lines 22). The latest committed data are updated with those in the write set.
@@ 4.6. Proof of Correctness
To prove that CCaaLF ensures serializability, we show that any transaction schedule allowed by CCaaLF can be transformed into a serial schedule in which the committed transactions are executed one by one in isolation. In particular, we show that the transaction execution results are the same when running with the CCaaLF schedule as when running with the serial schedule.
Definitions and Setup:. Let:
  • $T_{1},T_{2},\dots,T_{n}$ be a set of transactions in the system.
@@
  • A schedule $S$ is an ordered sequence of read/write operations from these transactions.
@@
  • A schedule is serial if all operations of a transaction $T_{i}$ execute without interleaving with operations from other transactions.
  • A schedule is serializable if it is equivalent to some serial schedule, i.e., it produces the same final state of the system.
@@
We aim to prove that the schedule $S$ produced by CCaaLF is serializable, i.e., there exists a serial schedule $S^{\prime}$ that is equivalent to $S$.
Transformation to Serial Schedule:.
  1. (1)
    Serialization Point: Denote the serialization point of a transaction $T_{i}$ by $t(T_{i})$. This is the logical time at which $T_{i}$ is considered to have committed in the global order of transactions. This point corresponds to the time when $T_{i}$ is validated in the algorithm (Algorithm 2, line 2).
  2. (2)
    Correctness Condition: We need to show that all operations from the schedule $S$ can be transformed into a serial schedule $S^{\prime}$, without changing the results of the transaction execution. We do this by shifting operations to their serialization points.
  3. (3)
    Read Operations: Let $r(x)$ denote a read operation on data item $x$ by transaction $T_{i}$.
    • Since read operations do not modify data, moving $r(x)$ to the serialization point $t(T_{i})$ does not change the final state of any data objects.
    • Moreover, the validation step (lines 22) ensures that the value read by $r(x)$ is consistent. Specifically, it guarantees that $x$ was not modified between the read and the serialization point.
    • Hence, moving $r(x)$ to $t(T_{i})$ preserves the correctness of the read.
  4. (4)
    Write Operations: Let $w(x)$ denote a write operation on data item $x$ by transaction $T_{i}$.
    • Write operations are protected by locks (lines 22). Therefore, their order is preserved with respect to other conflicting writes, ensuring that moving $w(x)$ to $t(T_{i})$ does not change the final state of any data objects.
    • Since read operations are validated against the committed state (ensured by lines 22), no read operation $r(x)$ from $T_{j}$ can access a value written by a transaction $T_{i}$ unless $T_{i}$’s serialization point precedes the validation of $T_{j}$.
    • In this case, the successful validation also indicates that no other write on $x$ exists in $S$ from $T_{i}$’s serialization point $t(T_{i})$ to $T_{j}$’s $r(x)$.
    • Therefore, moving $w(x)$ to $t(T_{i})$ does not affect any reads in $S$.
Since we can shift all operations in $S$ to their respective serialization points without altering the execution results, the final state of the system under $S$ is identical to that of a serial schedule $S^{\prime}$, where transactions are executed one-by-one according to their serialization points. Thus, every schedule $S$ allowed by CCaaLF is serializable. This completes the proof that CCaaLF ensures serializability.
5. Learning Agent Function
In this section, we first define the optimization of the agent function as a black-box optimization problem. We then describe how CCaaLF efficiently searches for the optimal agent functions.
5.1. Problem Definition
CCaaLF starts the optimization process for the agent function when it detects performance degradation caused by workload drifts. It uses throughput as the performance metrics, similar to previous works (DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/cidr/TangJE17/ACC, ).
Let $\mathbf{S}$ be the set of possible input states $\mathbf{{s}}$, and $\mathbf{A}$ be the set of possible output actions $\mathbf{{a}}$. The optimization problem is defined as:
(6) maximize $\text{Score}(\mathcal{{F}})$
@@ 5.1. Problem Definition
We observe that different actions are amenable to different optimization techniques. Model-based optimizations are suitable for some actions, whereas search-based techniques are suitable for others. Furthermore, different actions have different impacts on the overall throughput, and thus the order in which they are optimized can affect convergence speed. Based on these observations, we decompose $\mathcal{{F}}$ into three sub-functions: $\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, and $\mathcal{{F}}_{P}$, and optimize them separately. Their outputs are combined to obtain $\mathbf{{a}}$ $=\langle{\mathbf{{a}}}_{d},{\mathbf{{a}}}_{t},{\mathbf{{a}}}_{p},({w_{1}},\dots,{w_{n}},{expose})\rangle$. $\mathcal{{F}}_{T}$ controls the timeout values, corresponding to ${\mathbf{{a}}}_{t}$. $\mathcal{{F}}_{D}$ controls the conflict detection and priority, corresponding to (${\mathbf{{a}}}_{d},{\mathbf{{a}}}_{p}$). $\mathcal{{F}}_{P}$ controls pipeline-wait actions, corresponding to $({w_{1}},\dots,{w_{n}},{expose})$.
@@
Learning the optimal agent function requires evaluating the value of $\mathcal{{F}}$). Existing works (DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/cidr/TangJE17/ACC, ) load the function into the running database and record the system throughput as the performance scores. This approach is time-consuming. In particular, the process of loading and running the system to evaluate $\mathcal{{F}}$) takes seconds  [^1] This process takes 3.68 seconds on average under our default experimental setting, as detailed in Section 6.1: TPCC workload, 16 threads, one warehouse. to complete and is repeated at every optimization step, adding significant overheads to the learning process (DBLP:conf/osdi/WangDWCW0021/polyjuice, ). In addition, the function $\mathcal{{F}}$) is non-smooth due to its discrete input $\mathcal{{F}}_{P}$. This non-smoothness makes it difficult to optimize $\mathcal{{F}}_{P}$ with model-based optimization methods. The following subsections discuss how CCaaLF addresses the aforementioned limitations of the existing approaches for evaluating $\mathcal{{F}}$) in two ways. First, it employs an online-trained surrogate model to minimize the number of performance score evaluations (Section 5.2). Second, it uses a graph reduction algorithm to effectively optimize $\mathcal{{F}}_{P}$ (Section 5.3).
5.2. Agent Function Learning with Online-Trained Surrogate Model
@@
As $\mathcal{{F}}$) is costly to evaluate, CCaaLF treats it as a black-box (regis2013combining/DYCORS, ; DBLP:conf/kdd/GolovinSMKKS17/Vizier, ; DBLP:books/daglib/0022247, ) and adopts a Bayesian optimization approach. The Bayesian optimization can efficiently identify high-performing agent functions with minimal evaluations, thereby reducing the evaluation cost. In particular, CCaaLF employs a probabilistic surrogate model, i.e., the Gaussian process model, which is trained online to approximate $\mathcal{{F}}$). The Gaussian process model is differentiable for efficient optimization and is well-suited as a surrogate model because it predicts both the mean and variance of performance scores, balancing exploration of new areas in the function space with exploitation of known high-performing areas.
@@
This surrogate model guides the optimization process by identifying the most promising agent function, i.e., the one expected to yield the highest performance score improvement, for subsequent deployment and evaluation. Specifically, during an optimization step, the surrogate model is first fitted to the historical data collected since the last workload drift, which optimizes its mean vector and covariance matrix to maximize the data likelihood. Next, for an unevaluated agent function $\mathcal{{F}}$, the fitted model predicts the mean $\mu(\text{Score}(\mathcal{{F}}))$ and variance $\sigma(\text{Score}(\mathcal{{F}}))$ of its performance score. After that, the Upper Confidence Bound (UCB) acquisition function computes a score for $\mathcal{{F}}$ as $\text{UCB}(\mathcal{{F}})=\mu(\text{Score}(\mathcal{{F}}))+\lambda\sigma(\text{Score}(\mathcal{{F}}))$, where $\lambda$ is a trade-off parameter and is set to $2.576$ in experiments, corresponding to the 99th percentile of a standard normal distribution to achieve an effective balance between exploration and exploitation. Note that the UCB score is differentiable with respect to the parameters of $\mathcal{{F}}$, and therefore CCaaLF can explicitly optimize these parameters to maximize the UCB score using the L-BFGS-B (DBLP:journals/mp/LiuN89, ) algorithm, which produces $\mathcal{{F}}_{next}$, the agent function with the highest expected performance score. Once $\mathcal{{F}}_{next}$ is identified, it is evaluated in the database system, and its actual performance score is recorded. Finally, the new data point of $\mathcal{{F}}_{next}$ and its score are added to the historical data, and the surrogate model is updated incrementally. This process is repeated iteratively, progressively refining predictions of $\mathcal{{F}}$) and identifying better agent functions over time.

### v4

@@ 9     $S\leftarrow$ all operations from conflict transactions
14          bar $\leftarrow$ $w_{T^{\prime}.type}$
15           $S\leftarrow S\cup\{\text{the first bar-th accesses in }T^{\prime}\}$
16
17    else if EarlyValidation(T)(T) = Abort then
@@ 24 $result\leftarrow$ SafeExecute($T,op$)
27     if EarlyValidation(T)(T) = Abort then
@@ 4.4.1. Execution
A transaction in NeurCC has the following fields:
@@
For each operation $op$, NeurCC maintains the following information:
@@
  • $op.key$: the unique key of the data object accessed by $op$.
@@
  • $op.priority$: the priority of $op$ when resolving its conflicts with other operations.
@@
Algorithm 1 describes how NeurCC executes an operation. It performs early validation to proactively abort transactions that will fail later (line  1). It also performs early validation before exposing a write (line 1), where the failure of a transaction $T$ could lead to the cascading abort of all transactions depending on it. By performing early validation (Algorithm 2), it reduces the impact of cascading aborts. The algorithm does not perform early validation for “detect all” because doing so incurs additional overhead and can restrict the search space. For example, it would prevent NeurCC from replicating 2PL, which does not incur the overhead of early validation checks.
NeurCC does not outperform 2PL when all the actions are detect-all, nor does it introduce a new lock-based protocol. However, we note that despite the extra validation cost compared to classical algorithms such as 2PL, the performance gained from the learned concurrency control actions outweighs this cost and leads to higher throughput, as will be illustrated in Section 7. Furthermore, when all actions in the agent function are detect all, NeurCC can skip validation mechanisms, including the tracking of $T.rs,T.ws,T.validate$ and the transaction validation at commit time, to further reduce the overhead over 2PL.
The algorithm distinguishes between stored procedure and interactive transaction modes. In the former, all the data accesses are known prior to transaction execution. In the latter, they are not known before transaction execution. For interactive transactions, the algorithm disallows actions that could violate correctness. Such actions include dirty reads and partial retry, which are disallowed because the behavior of the conflicting transaction is unknown, and because the user has already seen the outputs of some operations.
@@
Given an operator $op$, the algorithm first invokes $GetCC(T,op)$ to get the action $\mathbf{{a}}$ (line 1). Next, it executes conflict detection and stores the conflicts in $S$. This step is skipped if ${\mathbf{{a}}}_{d}$ is no detection, and the algorithm enforces $op$ to read clean data (lines 11). If ${\mathbf{{a}}}_{d}$ is detect critical, it performs early validation to check if previously read data have changed (line 1) and aborts $T$ if the validation fails. The conflicts with lower priorities than that of $op$ are removed from $S$. The remaining are resolved with the given timeout ${\mathbf{{a}}}_{t}$ (lines 11). If not all the conflicts in $S$ are resolved within ${\mathbf{{a}}}_{t}$, $T$ is aborted (line 1). Otherwise, the algorithm proceeds to read the latest committed data (line 1), or to write to its write set $T.ws$ (line 1).
@@
For stored procedure transactions, the algorithm allows dirty reads. In particular, it uses $w_{1},\dots,w_{n}$ to detect critical conflicts before dirty reads, uses exposeexpose to decide whether to make uncommitted write visible to other transactions, and records dependencies caused by dirty reads in $T.dep$ (lines 1 and 1). Each data object has a version chain containing dirty data. The execution when ${\mathbf{{a}}}_{d}$ is no detection or detect all is the same as that of interactive transactions. When ${\mathbf{{a}}}_{d}$ is detect critical, it detects critical conflicts between $T$ and its dependent transactions in $T.dep$ (lines  11). In particular, for every $T^{\prime}\in T.dep$, it adds the first $w_{T^{\prime}.type}$ accesses of $T^{\prime}$ to $S$. Once all the conflicts in $S$ are resolved, it lets $op$ read the data at the tail of the corresponding version chain, and updates $T.dep$ to track the read-write dependency (lines 11). After executing $op$ (line 1), if $expose=1$, it performs early validation and aborts $T$ if the validation fails. If the validation succeeds, it exposes the uncommitted data in $T.ws$ by appending them to the corresponding version chains.
We use the ASOCC example to illustrate how a transaction is executed with NeurCC. We assume the transaction is in the interactive transaction mode. Before executing an operation, NeurCC first computes the input state $\mathbf{{s}}$ and retrieves the corresponding actions by evaluating $\mathcal{{F}}(\mathbf{{s}})$ (line 1). If the selected action is no detection (i.e., the accessed data is cold), NeurCC executes it optimistically without blocking (line 1). If the action is detect all (i.e., the accessed data is hot), NeurCC performs conflict detection before execution (line 1), and upon detecting a conflict, it blocks or abort the transaction according to the selected timeout (line 1). If the action is detect critical (i.e., the accessed data is warm), NeurCC performs early validation (line 1), and if validation fails, it aborts the transaction immediately.
1Function EarlyValidation($T$):
@@ 9 Function Commit($T$):
18     if $op.tx\neq op^{\prime}.tx$ or $op$ is locked by other transactions then return Abort
19
@@
22     Append opop to tuple($op.key$)’s version chain as clean data
@@ 4.4.2. Validation and commit
Algorithm 2 shows how the algorithm performs an OCC-like four-step validation at commit time. First, it waits for all dependent transactions in $T.dep$ to complete, and aborts if any dependent transaction aborts (lines 22). Second, it locks the private write set to avoid concurrent read of the data in the set (line 2). Third, it verifies that data in the read set are the same as the latest committed versions (lines 22). It aborts if the verification fails. Fourth, it removes dirty data from $T$ on the corresponding version chains and unlocks the write set (lines 22). The latest committed data is updated to those in the write set.
@@ 4.5. Proof of Correctness
To prove that NeurCC ensures serializability, we show that any transaction schedule allowed by NeurCC can be transformed into a serial schedule in which the committed transactions are executed one by one in isolation. In particular, we show that the transaction execution results are the same when running with the NeurCC schedule as when running with the serial schedule.
Definitions and Setup. Let:
  • Each transaction $T$ is assigned a unique timestamp $t(T)$ at its serialization point (Algorithm 2, line 2).
  • Let $T_{1},T_{2},\dots,T_{n}$ be the set of committed transactions, indexed such that $t(T_{1})<t(T_{2})<\cdots<t(T_{n})$, where $t(T_{i})$ denotes the serialization point of transaction $T_{i}$.
@@
  • A schedule $S$ is an interleaved sequence of read and write operations from $T_{1},T_{2},\dots,T_{n}$.
  • Two schedules of committed transactions are equivalent if, for every committed transaction, all of its read operations return the same values in both schedules, and the two schedules result in the same final database state.
@@
  • A schedule is serial if all operations of each transaction execute contiguously, without interleaving with other transactions.
  • A schedule is serializable if it is equivalent to some serial schedule.
  • Let schedule $S_{k}$ be a subsequence of $S$ containing all operations from transactions $T_{1},T_{2},\dots,T_{k}$ ($S_{n}=S$).
  • Let schedule $S^{\prime}_{k}$ be a serial schedule $\langle T_{1},T_{2},\dots,T_{k}\rangle$ in which $T_{1},T_{2},\dots T_{k}$ are executed sequentially.
@@
We aim to prove that any schedule $S$ produced by NeurCC is serializable. That is, $S_{n}$ is equivalent to a serial schedule $S^{\prime}_{n}$ in which committed transactions $T_{1},T_{2},\dots,T_{n}$ are executed sequentially. We prove by induction on $k$ (from $1$ to $n$) that $S_{k}$ is equivalent to $S^{\prime}_{k}$.
Base Case (k = 1). For a single committed transaction $T_{1}$, the schedule $S_{1}$ trivially corresponds to the serial schedule $\langle T_{1}\rangle$, as there is no transaction operation interleaving.
Inductive Step. Assume that for $T_{1},T_{2},\dots,T_{k}$, the schedule $S_{k}$ is equivalent to the serial schedule $S^{\prime}_{k}$. Now consider an additional committed transaction $T_{k+1}$, with $t(T_{k})<t(T_{k+1})$. We show that inserting operations from $T_{k+1}$ to $S_{k}$ to form $S_{k+1}$ still produces the same return values and final state as the schedule $S^{\prime}_{k+1}$.
Let $r(x)$ denote a read operation on data object $x$ from transaction $T_{k+1}$, $w(x)$ denote a write operation on data object $x$ from transaction $T_{k+1}$. We now prove that $S_{k+1}$ and $S^{\prime}_{k+1}$ produce the same return values and final database state.
  • $T_{1},\dots,T_{k}$ read the same return values in $S_{k+1}$ and $S^{\prime}_{k+1}$: For any transaction $T_{i}\in\{T_{1},\dots,T_{k}\}$, all of its read operations occur before its serialization point $t(T_{i})$. Since any write $w(x)\in T_{k+1}$ becomes visible only after $t(T_{k+1})$, and $t(T_{i})<t(T_{k+1})$, none of the reads in $T_{i}$ can observe writes from $T_{k+1}$. Furthermore, since $T_{1},\dots,T_{k}$ read the same return values in $S_{k}$ and $S^{\prime}_{k}$, and the state visible to them remains unchanged considering updates from $T_{k+1}$, they continue to read the same values in $S_{k+1}$ and $S^{\prime}_{k+1}$.
  • $T_{k+1}$ reads the same return values in $S_{k+1}$ and $S^{\prime}_{k+1}$: Since $t(T_{k+1})>t(T_{i})$ for all $i\in\{1,2,\dots,k\}$, any read operation $r(x)\in T_{k+1}$ occurs after the updates to $x$ from $S_{k}$. Otherwise, $T_{k+1}$ would detect a conflicting write to $x$ from a transaction $T_{i}$ with $t(T_{i})<t(T_{k+1})$, and abort due to a failed validation of its read set (Algorithm 2, line 2). In summary, each $r(x)\in T_{k+1}$ observes the final state of $x$ resulting from all updates in $S_{k}$. Since the final database states produced by $S_{k}$ and $S^{\prime}_{k}$ are identical, and $T_{k+1}$ reads from this final state in both $S_{k+1}$ and $S^{\prime}_{k+1}$, $T_{k+1}$ reads the same return values in $S_{k+1}$ and $S^{\prime}_{k+1}$.
  • $S_{k+1}$ and $S^{\prime}_{k+1}$ produce the same final database state: Since $t(T_{k+1})>t(T_{i})$ for all $i\in\{1,2,\dots,k\}$, any update operation $w(x)\in T_{k+1}$ occurs after the updates to $x$ from $S_{k}$. Otherwise, $w(x)$ would have been blocked during write lock acquisition (Algorithm 2, line 2). This guarantees that no update in $S_{k}$ overwrites or interferes with the updates made by $T_{k+1}$, ensuring that $S_{k+1}$ produces the same final state as $S^{\prime}_{k+1}$ for all $x$ modified by $T_{k+1}$. For data objects not modified by $T_{k+1}$, the final state is already identical in $S_{k}$ and $S^{\prime}_{k}$, so the resulting state remains the same in both $S_{k+1}$ and $S^{\prime}_{k+1}$.
By induction, for all $k\in[1,n]$, we have $S_{k}$ equivalent to $S^{\prime}_{k}$. Hence, the full schedule $S=S_{n}$ is equivalent to the serial schedule $S^{\prime}_{n}$. Therefore, NeurCC ensures serializability.
5. Learning Function and Feature Selector
@@
In this section, we describe how NeurCC learns the function $\mathcal{{F}}$. We observe that different actions are amenable to different optimization techniques. More specifically, model-based optimizations are suitable for some actions, whereas search-based techniques are suitable for others. Furthermore, different actions have different impacts on the overall throughput, and thus the order in which they are optimized can affect convergence speed. Based on these observations, we decompose $\mathcal{{F}}$ into three sub-functions: $\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, and $\mathcal{{F}}_{P}$, and optimize them separately. Their outputs are combined to obtain $\mathbf{{a}}$ $\langle{\mathbf{{a}}}_{d},{\mathbf{{a}}}_{t},{\mathbf{{a}}}_{p},({w_{1}},\dots,{w_{n}},{expose})\rangle$. $\mathcal{{F}}_{T}$ controls the timeout values, corresponding to ${\mathbf{{a}}}_{t}$. $\mathcal{{F}}_{D}$ controls the conflict detection and priority, corresponding to (${\mathbf{{a}}}_{d},{\mathbf{{a}}}_{p}$). $\mathcal{{F}}_{P}$ controls pipeline-wait actions, corresponding to $({w_{1}},\dots,{w_{n}},{expose})$.
@@
Learning the optimal function requires evaluating the performance of $\mathcal{{F}}$, i.e., Score($\mathcal{{F}}$). Existing works (Tang and Elmore, 2018; Wang et al., 2021; Tang et al., 2017) load the function into the running database and use the system throughput as Score($\mathcal{{F}}$). This approach is time-consuming, because loading and running the system takes seconds [^2] This process takes 3.68 seconds on average under our default experimental setting, as detailed in Section 7.1: TPCC workload, 16 threads, one warehouse. to complete, and they need to be repeated at every optimization step, adding significant overheads to the learning process (Wang et al., 2021). In addition, Score($\mathcal{{F}}$) is non-smooth due to its discrete input $\mathcal{{F}}_{P}$, which makes it difficult to optimize $\mathcal{{F}}_{P}$ with model-based optimization methods. NeurCC addresses the aforementioned limitations of the existing approaches for evaluating Score($\mathcal{{F}}$) in two ways. First, it employs an online-trained surrogate model to minimize the number of performance evaluations (Section 5.1). Second, it uses a graph reduction search algorithm to effectively optimize $\mathcal{{F}}_{P}$ (Section 5.2).
5.1. Learning Function Via Surrogate Model
@@
As Score($\mathcal{{F}}$) is costly to evaluate, NeurCC treats it as a black-box (Regis and Shoemaker, 2013; Golovin et al., 2017; Conn et al., 2009) and adopts a Bayesian optimization approach. Bayesian optimization can efficiently identify high-performing functions with a small number of evaluations, thereby reducing the cost. In particular, NeurCC employs a probabilistic surrogate model, i.e., the Gaussian process model, which is trained online to approximate Score($\mathcal{{F}}$). This model is differentiable for efficient optimization and is well-suited as a surrogate model because it predicts both the mean and variance of the performance, which balances exploration of new areas in the function space with exploitation of known high-performing areas.
@@
The surrogate model guides the optimization process by identifying the most promising function, i.e., the one expected to yield the highest performance improvement, for subsequent deployment and evaluation. Specifically, during an optimization step, the surrogate model is first fitted to the historical data collected since the last workload drift, which optimizes its mean vector and covariance matrix to maximize the data likelihood. Next, for an unevaluated function $\mathcal{{F}}$, the fitted model predicts the mean $\mu(\text{Score}(\mathcal{{F}}))$ and variance $\sigma(\text{Score}(\mathcal{{F}}))$ of its performance. After that, the Upper Confidence Bound (UCB) acquisition function computes a score for $\mathcal{{F}}$ as $\text{UCB}(\mathcal{{F}})=\mu(\text{Score}(\mathcal{{F}}))+\lambda\sigma(\text{Score}(\mathcal{{F}}))$, where $\lambda$ is a trade-off parameter. It is set to $2.576$ in our experiments, corresponding to the 99th percentile of a standard normal distribution to achieve an effective balance between exploration and exploitation. Note that the UCB score is differentiable with respect to the parameters of $\mathcal{{F}}$, and therefore NeurCC can explicitly optimize these parameters to maximize the UCB score using the L-BFGS-B (Liu and Nocedal, 1989) algorithm, which produces $\mathcal{{F}}_{next}$, the function with the highest expected performance. Once $\mathcal{{F}}_{next}$ is identified, it is evaluated in the database system, and its actual performance score is recorded. Finally, the new data point of $\mathcal{{F}}_{next}$ and its score are added to the historical data, and the surrogate model is updated. This process is repeated, progressively refining predictions of Score($\mathcal{{F}}$) and identifying better functions over time.